OpenStudy (anonymous):

How long will it take for $2000 to double if it is invested at 6.25% interest compounded continuously? can some show me how to set this problem into this equation A=Pe^rt, please!

OpenStudy (mertsj):

\[4000=2000e ^{.0625t}\]

OpenStudy (mertsj):

\[2=e ^{.0625t}\]

OpenStudy (mertsj):

\[\ln 2=.0625t(\ln e)\]

OpenStudy (mertsj):

\[\frac{\ln2}{.0625}=t\]

OpenStudy (mertsj):

\[11.09 yrs = t\]

OpenStudy (anonymous):

can you please give me a crash course on what you did, Please...

OpenStudy (mertsj):

Did you understand how the values were substituted into the equation?

OpenStudy (anonymous):

no i didnt

OpenStudy (mertsj):

A is the amount which was determined to be $4000 because the problem asked how long it would take a $2000 investment to double.

OpenStudy (mertsj):

If $2000 doubles, it will then be $4000 would you agree?

OpenStudy (mertsj):

So the problem is really asking "how long will it take $2000 to turn into $4000"

OpenStudy (anonymous):

what does the e stand for?

OpenStudy (mertsj):

The e is a frequently used mathematical constant. It doesn't stand for anything. It is sort of like pi. Just a commonly used constant. It is part of the formula just like pi is when you write

OpenStudy (mertsj):

\[A=\pi r^2\]

OpenStudy (mertsj):

You don't replace pi with anything and you don't replace e with anything.

OpenStudy (anonymous):

can you help me figure out another problem but let me do it and maybe you can coach me.

OpenStudy (mertsj):

It's numerical value (if you must know) is 2.718281828...

OpenStudy (mertsj):

ok

OpenStudy (mertsj):

Lay it on me.

OpenStudy (anonymous):

What rate of interest compounded continuously is needed for an investment of $500 to grow to $900 in 10 years?

OpenStudy (mertsj):

What is A?

OpenStudy (anonymous):

900

OpenStudy (mertsj):

What is P?

OpenStudy (anonymous):

500

OpenStudy (mertsj):

What is t?

OpenStudy (anonymous):

10

OpenStudy (mertsj):

What are you trying to find?

OpenStudy (anonymous):

the rate of interest

OpenStudy (mertsj):

Yes. r

OpenStudy (mertsj):

So plug everything in.

OpenStudy (anonymous):

900=500e^r10

OpenStudy (mertsj):

\[900=500e ^{10r}\]

OpenStudy (mertsj):

Divide both sides by 500

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

In1.8=e^10r

OpenStudy (mertsj):

\[1.8=e ^{10r}\]

OpenStudy (mertsj):

Now when you take the natural log of both sides it should look like this: \[\ln 1.8=\ln e ^{10r}\]

OpenStudy (anonymous):

ok

OpenStudy (mertsj):

Now bring the exponent down using the third law of logs and get this

OpenStudy (mertsj):

\[\ln 1.8=10r \ln e\]

OpenStudy (mertsj):

But of course ln e = 1 so we have:

OpenStudy (mertsj):

\[\ln 1.8=10r\]

OpenStudy (mertsj):

So just divide both sides by 10 and you are done

OpenStudy (anonymous):

0.18=r

OpenStudy (mertsj):

\[\frac{\ln 1.8}{10}=r=.05878=5.88 %\]

OpenStudy (mertsj):

Did you take the natural log of 1.8 before you divided by 10?

OpenStudy (anonymous):

no

OpenStudy (mertsj):

Do you now how to do that?

OpenStudy (anonymous):

no

OpenStudy (mertsj):

You do understand that it says ln 1.8 divided by 10?

OpenStudy (mertsj):

Do you have a scientific calculator?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

Can you find the ln button?

OpenStudy (mertsj):

It's probably right next to the log button

OpenStudy (anonymous):

yes i just pressed it and got the same as you did do i now dive that by 10

OpenStudy (mertsj):

Yes

OpenStudy (mertsj):

Now multiply by 100 to change it to percent.

OpenStudy (anonymous):

got it you just answered my question!! thank you so much Mertsj

OpenStudy (mertsj):

You're welcome and good luck. Isn't learning fun???!!!

OpenStudy (anonymous):

yeah your the best!!!!!

OpenStudy (mertsj):

Thanks.